allycat:

The price of a technology stock has risen to $9.77 today. Yesterday's price was $9.66 . Find the percentage increase. Round your answer to the nearest tenth of a percent.

Nnesha:

\[\large\rm \frac{ \left| today's~ price - yesterday's ~~price \right| }{ yesterday's ~price } \times 100\]